OpenStudy (anonymous):

The perimeter of right triangle RST is equal to the perimeter of isosceles triangle XYZ. The lengths of the legs of the right triangle are 6 and 8. If the length of each side of the isosceles triangle is an integer, what is the greatest possible length for one of the sides of the isosceles triangle XYZ?

OpenStudy (anonymous):

The lengths of the legs are 6, 8, by pythagoras the hypotenuse is 10 for a total of 24

OpenStudy (anonymous):

Is the pythagoras theorm \[a ^{2}+b ^{2}\]

OpenStudy (anonymous):

\(a^2+b^2=c^2\) and so \[6^2+8^2=c^2\] tells you \(c=10\)

OpenStudy (anonymous):

So the type of triangle does not matter?

OpenStudy (anonymous):

as for the isosceles triangle, i guess the sides could be 11, 11, and 2 not sure if you can get a longer side

OpenStudy (anonymous):

no it matter that is for a right triangle

OpenStudy (anonymous):

For the isosceles triangle then it one of the side has to be 10? The problem had options

OpenStudy (wolf1728):

How about 22 1 and 1 It asks for length of longest side

OpenStudy (anonymous):

i think two sides could be 11, the third side 2

OpenStudy (anonymous):

A.) 10 B.) 11 C.) 14 D.) 16 E.) 22

OpenStudy (wolf1728):

I'd say E

OpenStudy (anonymous):

there is no triangle with two sides 1, and one side 22

OpenStudy (anonymous):

try to draw it and you will see why

OpenStudy (anonymous):

but there is a 11, 11, 2, triangle

OpenStudy (wolf1728):

Geez yes - I was interested in the mathematical solution. I neglected the geometry.

OpenStudy (anonymous):

Geometry is all logic haha

OpenStudy (wolf1728):

Yes the Triangle Inequality Theorem Any side of a triangle must be: less than the sum of the other 2 sides and greater than the difference of the other 2 sides.

OpenStudy (anonymous):

So which answer is best 10 or 11?

OpenStudy (wolf1728):

I'd say satellite73 had the correct solution 11

OpenStudy (anonymous):

Alright, Thank You Both! c:

OpenStudy (wolf1728):

u r welcome (and I'd better slow down in my anxiety to find the "correct" solution)